Dominique MANCHON UMR 6620, 24 av. des Landais, 63177 Aubire cedex

Tel : 04 73 40 76 98. Adresse électronique : manchon@math.univ-bpclermont.fr

Liste des publications parues ou à paraître

(Janvier 2012)

Revues avec comité de lecture :

[1] Formule de Weyl pour les groupes de Lie nilpotents, J. f.d. Reine u. Angew. Math. 418, 77-129 (1991).

[2] Calcul symbolique sur les groupes de Lie nilpotents et applications, J. Funct. Anal. 102 (2), 206-251 (1991).

[3] Weyl symbolic calculus on any Lie group, Acta Appl. Math. 30, 159-186 (1993).

[4] Opérateurs pseudo-differentiels et représentations unitaires des groupes de Lie, Bull. Soc. Math. France 123, 117-138 (1995).

[5] Opérateurs aux différences finies, calcul pseudo-différentiel et représentations des groupes de Lie (avec M. Andler), J. Geom. Phys. 27, 1-29 (1998).

[6] L’algèbre de Hopf bitensorielle, Comm. Alg. 25 (5), 1537-1551 (1997).

[7] Distributions à support compact et représentations unitaires, J. Lie Theory 9, 403-424 (1999).

[8] Front d’onde et propagation des singularités pour un vecteur-distribution, Coll. Math. 81 (2), 161-191 (1999).

[9] Une remarque sur l’exponentielle-etoile, Rencontres math. de Glanon (1997).

[10] Choix des signes pour la formalité de M. Kontsevich (avec D. Arnal et M. Masmoudi), Pacific J. Math. 203, N1, 23-66 (2002).

[11] Poisson bracket, deformed bracket and gauge group actions in Kontsevich deformation quantization, Lett. Math. Phys. 52, 301-310 (2000).

[12] On quantization of quadratic Poisson structures (avec M. Masmoudi et A. Roux), Commun. Math. Phys. 225, 121-130 (2002).

[13] Cohomologie tangente et cup-produit pour la quantification de Kontsevich (avec Ch. Torossian), Annales de l’Université Blaise Pascal 10, 75-106 (2003).

[14] Orbites coadjointes et variétés caractéristiques (avec A. Baklouti et S. Dhieb), J. Geom. Phys. 54, 1-41 (2005).

[16] Shuffle relations for regularised integrals of symbols (avec S. Paycha), Comm. Math. Phys. 270, 13-51 (2007).

[17] Birkhoff type decompositions and the Baker-Campbell-Hausdorff recursion (avec K. Ebrahimi-Fard et L. Guo), Comm. Math. Phys. 267, 821-845 (2006).

[18] On Matrix Differential Equations in the Hopf Algebra of Renormalization (avec K. Ebrahimi-Fard), Adv. Theor. Math. Phys. 10, 879-913 (2006).

[21] New identities in dendriform algebras, (avec K. Ebrahimi-Fard et F. Patras), Journal of Algebra 320, 708-727 (2008).

[22] A Magnus- and Fer-type formula in dendriform algebras, (avec K. Ebrahimi-Fard), Found. of Comput. Math. Volume 9, Number 3, 295-316 (2009).

[23] A noncommutative Bohnenblust-Spitzer identity for Rota-Baxter algebras solves Bogoliubov’s recursion (avec K. Ebrahimi-Fard et F. Patras), J. Noncommutative Geometry Vol. 3, Issue 2, 181-222 (2009).

[25] Confluence of singularities of a differential equation: a Lie algebra contraction approach (avec M. B. Zahaf), Int. J. of Math. Analysis, Online Edition, Vol. 3 No. 1-4, 23-40 (2009).

[26] Quelques aspects combinatoires de la renormalisation, Gazette des Mathématiciens No 119, 5-15 (2009).

[27] Dendriform equations (avec K. Ebrahimi-Fard), Journal of Algebra 322, 4053-4079 (2009).

[28] Two interacting Hopf algebras of trees (avec D. Calaque et K. Ebrahimi-Fard), Advances in Applied Mathematics 47, no. 2, 282-308 (2011).

[31] Nested sums of symbols and renormalised multiple zeta functions (avec S. Paycha), Int. Math. Res. Notices 2010 issue 24, 4628-4697 (2010).

[33] Lois pré-Lie en interaction (avec A. Saïdi), Comm. Alg. volume 39 No 10, 3662-3680 (2011).

Actes de congrès avec comité de lecture :

[15] Bogota lectures on Hopf algebras, from basics to applications to renormalization, Comptesrendus des Rencontres mathématiques de Glanon 2001 (parus en 2003) et arXiv:math. QA/0408405.

[24] The combinatorics of Bogoliubov’s recursion in renormalization (avec K. Ebrahimi-Fard), CIRM 2006 workshop ”Renormalization and Galois Theories”, IRMA Lect. in Math. and Theor. Phys. 15 (A. Connes, F. Fauvet, J.-P. Ramis eds.), 179-207 (2009).

[29] Connected filtered Hopf algebras and renormalization, in “Motives, Quantum Field Theory and pseudodifferential Operators”Clay Math. Proceedings vol. 12. A. Carey, D. Ellwood, S. Paycha, S. Rosenberg Eds. (2010).

[30] Renormalised multiple zeta values which respect quasi-shuffle relations, in ”Combinatorics and physics”, K. Ebrahimi-Fard, M. Marcolli, W. Van Suijlekom eds., Contemp. Math. No 539 (2011).

[36] A deformation approach of the Kirillov map for exponential lie groups (avec A. Baklouti et S. Dhieb), Congrès tuniso-japonais de Kerkennah (nov. 2009), Adv. in Pure and Appl. Math. 2, 421-436 (2011).

[35] A short survey on pre-Lie algebras, E. Schrödinger Institut Lectures in Mathematics and Physics, A. Carey Ed., Eur. math. Soc. (2011).

Chapitres de livres :

[19] Hopf algebras in renormalisation, Encycl. Math. (note brève , à paraître).

[20] Hopf algebras and renormalisation, Handbook of Algebra, Vol. 5 (M. Hazewinkel ed.), 365-427 (2008) (version augmentée et actualisée de [15]).